OpenStudy (anonymous):

A glass tubing is 0.525 m long. How many inches long is the tubing

OpenStudy (tkhunny):

\(0.525\;m\cdot\dfrac{100\;cm}{1\;m}\cdot\dfrac{1\;in}{2.54\;cm}\) Look at each piece until you see what it is.